Obtain the solution of the differential equation


y"+3y'+2y=e^(2t) using the method of variation of parameters

1
Expert's answer
2022-03-29T06:53:35-0400

"y''+3y'+2y=e^{2t}\\\\Homogeneous\\,\\,equation:\\\\y''+3y'+2y=0\\\\\\lambda ^2+3\\lambda +2=0\\\\\\lambda _1=-2,\\lambda _2=-1\\\\y=C_1e^{-2t}+C_2e^{-t}\\\\The\\,\\,fundamental\\,\\,system\\,\\,of\\,\\,solutions:\\\\y_1=e^{-t},y_2=e^{-2t}\\\\y\\left( t \\right) =C_1\\left( t \\right) y_1+C_2\\left( t \\right) y_2\\\\By\\,\\,the\\,\\,method:\\\\\\left\\{ \\begin{array}{c}\ty_1\\left( t \\right) C_1'\\left( t \\right) +y_2\\left( t \\right) C_2'\\left( t \\right) =0\\\\\ty_1'\\left( t \\right) C_1'\\left( t \\right) +y_2'\\left( t \\right) C_2'\\left( t \\right) =e^{2t}\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\te^{-t}C_1'\\left( t \\right) +e^{-2t}C_2'\\left( t \\right) =0\\\\\t-e^{-t}C_1'\\left( t \\right) -2e^{-2t}C_2'\\left( t \\right) =e^{2t}\\\\\\end{array} \\right. \\Rightarrow \\\\\\Rightarrow \\left\\{ \\begin{array}{c}\tC_2'\\left( t \\right) =-e^tC_1'\\left( t \\right)\\\\\te^{-t}C_1'\\left( t \\right) =e^{2t}\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tC_2'\\left( t \\right) =-e^tC_1'\\left( t \\right)\\\\\tC_1'\\left( t \\right) =e^{3t}\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tC_2'\\left( t \\right) =-e^{4t}\\\\\tC_1'\\left( t \\right) =e^{3t}\\\\\\end{array} \\right. \\\\\\Rightarrow \\left\\{ \\begin{array}{c}\tC_1\\left( t \\right) =\\frac{1}{3}e^{3t}+C_1\\\\\tC_2\\left( t \\right) =-\\frac{1}{4}e^{4t}+C_2\\\\\\end{array} \\right. \\Rightarrow y\\left( t \\right) =\\left( \\frac{1}{3}e^{3t}+C_1 \\right) e^{-t}+\\left( -\\frac{1}{4}e^{4t}+C_2 \\right) e^{-2t}=\\\\=\\frac{1}{12}e^{2t}+C_1e^{-t}+C_2e^{-2t}"